OpenStudy (anonymous):

help! Elimination problems! need explaination! 4x + y=8.....-3x -y = 0

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

4x + y = 8 -3x - y = 0 add these two equations together... notice you'll end up with 1 equation with only 1 variable. this is why they call it elimination... you eliminated 1 variable... :)